OpenStudy (anonymous):

It takes a car 1.75 h to go from mile marker 10 to mile marker 115. What is the average speed of the car? A. 1.6 mph B. 60 mph C. 65 mph D. 183 mph

OpenStudy (anonymous):

b. 60 mph do you want me to explain the answer?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

and can you help with more

OpenStudy (anonymous):

A bicyclist is in a 50-km race. He says he had an average velocity of 35. What is missing in the cyclist’s velocity? A. units only B. units and distance C. direction only D. units and direction i think this one is c

OpenStudy (anonymous):

\[Average \ speed = \frac{Total \ Distance \ covered}{Total \ Time \ taken}\]

OpenStudy (anonymous):

From 10 mile marker to 115 mile marker, Distance is : \(115 - 10\) = \(105 \ miles\)

OpenStudy (anonymous):

Time taken is given to you ie \(1.75 \ hours\)

OpenStudy (anonymous):

\[Average \ speed = \frac{105 \ miles}{1.75 \ hours} = ??\]

OpenStudy (anonymous):

the top one is 60mph

OpenStudy (anonymous):

@Catlover5925

OpenStudy (anonymous):

In second, velocity has Magnitude as well as Direction.

OpenStudy (anonymous):

The units of velocity are : \(m/s\) or \(km/h\)

OpenStudy (anonymous):

units are not written there, alone 35 is meaningless.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Direction is also missing, so what should it be?

OpenStudy (anonymous):

d

OpenStudy (anonymous):

i have 2 more questions

OpenStudy (catlover5925):

im soooo sorry I don't know but the first question I think is D

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

i got 52 is that right

OpenStudy (anonymous):

@micahm

OpenStudy (micahm):

ok re post question

OpenStudy (anonymous):

Use the following table of a school bus during morning pickups to calculate its average speed between 2.8 km and 4.2 km. A. 42 km/h B. 50 km/h C. 52 km/h D. 55 km/h

OpenStudy (anonymous):

never,mind

OpenStudy (micahm):

60

OpenStudy (micahm):

B